OpenStudy (anonymous):

Write in general form the equation of the line that passes through the given points. 1. (4,-5) , (2,-5) 2. (9,4),(0,2)

OpenStudy (disqin):

Use the equation\[(y-y_1)=\frac{y_2-y_1}{x_2-x_1}(x-x_1)\] where x1,x2,y1 and y2 are your given points.

OpenStudy (anonymous):

40x-4y=20 is my answer?

OpenStudy (anonymous):

-4x+9y=0 in # 2?

OpenStudy (campbell_st):

here is a really simple method.... plots the points on the numberplane 

OpenStudy (campbell_st):

same for question 2  hope it helps